# iWISDM: Assessing instruction following in multimodal models at scale

### v1

@@ 4.2 Complexity
Using the AutoTask framework in iWISDM, we created three benchmarks corresponding to Low, Medium, and High complexity tasks. The Low complexity tasks were restricted to contain exactly one logical joining operator (”and/or”), exclude Switch operators, and require only boolean actions. Medium complexity tasks had the same logical joining operator restriction as low complexity and boolean action requirements, with the complexity increase coming from adding a Switch operator. Finally, high complexity tasks have between one and two logical joining operators, an included Switch operator, and require boolean and object property action responses (i.e. ’… category of object 1?’, Answer: ’planes’). For further details refer to Figure A1. Additionally, since the benchmarks are generated based on a synthetic stimuli dataset, while all LMMs are trained with naturalistic stimuli, we aim to confirm whether LMMs can recognize rendered ShapeNet objects. Therefore, we developed two sets of simplistic single-frame tasks: location-only and category-only existence tasks (e.g. Is object 1 in the bottom left?). Evaluation on these two single-frame task sets provides an upper bound of LMMs’ performance.
The set of tested models was limited due to the scarcity of applicable models. Many popular open-source LMMs, such as MiniGPT4 and LLaVa, were unable to perform the task simply due to their limited image sequence lengths. A minimum image sequence length of ten is needed to complete all complexity levels. We were able to properly evaluate two open source models, InternLM-XComposer2-7b and MMICL-Instructblip-T5-xl. For samples of prompts used to evaluate each model see Appendix A.2.
@@ 4.3 Results
Figure 4a shows the accuracy of actions taken by the models plotted against the complexity level for each type of prompt. GPT-4V generally achieves the best model scores, with the largest performance gap on the low and high complexity tasks. However, relative to human performance these gaps are marginal at best. Broadly, MMICL was the worst performing model. The expected inverse correlation between complexity and action accuracy was only captured clearly by the GPT-4V and Gemini-Pro-1.0 results.
@@
We also analyzed how each model performed on subsets of tasks where single or multiple object properties were involved. Figure 4 b-e shows the average accuracy of all models conditioned on task properties and complexity level. We did not include the object-identity subgroup in the high complexity plot of Figure 4e as non-boolean response types for object identity tasks are not feasible. The performances on high complexity tasks show a clear ranking between the abilities of GPT-4V and Gemini-pro models on tasks with diverse response options. For almost all models, and complexities, location-only tasks posed the most difficulties. This finding confirms previous analyses of GPT-4V, which found that it often struggles to correctly recognize an object’s position within an image [^1] https://blog.roboflow.com/gpt-4v-object-detection (Majumdar et al., 2024).
To further investigate the response patterns of the LMMs we performed further analyses on a subset of the models seen in Appendix Figures A7, A8, A9, and A10. To determine the effects of delay frames on model performance a set of simple delayed-match-to-sample tasks were generated which only differed in difficulty by the number of delay frames. Appendix Figure A7 shows that for InternLM-XComposer2, MMICL, and GPT-4v, the addition of single or multiple delay frames in a task has little effect on task performance. Next we looked into how the number of different boolean operators affected open-source model (InternLM-XComposer2 & MMICL) performance. The Appendix Figure A8 a-d results show that, generally, across all boolean operators an increase in their abundance within a task leads to worse model performance, which is expected. In Appendix Figure A9 we examined how the number of stimuli effected task performance across complexities. We expected to see that as the number of stimuli increases, the task performance would decrease. This was found to be the case for Low complexity tasks, which can be seen in Appendix Figure A9 a. However Medium and High complexity tasks displayed an inverse trend, seen in Appendix Figure A10 b and c. Finally, we investigated the exact effect that different required response types had on accuracy for the High complexity benchmark. As Appendix Figure A10 shows, when tasks required a response that was a non-boolean type word, the models performed significantly worse.
@@ 5 Conclusion and future directions
We introduced iWISDM as a platform for validating multimodal models. As a benchmark, our primary focus is on assessing the ability of LMMs to follow instructions in visual-language decision making tasks. We believe iWISDM would be an important benchmark that complements existing benchmarks which evaluate LMM capabilities in areas such as commonsense reasoning, numerical computation, or relational inferences (Fu et al., 2023). Although the current version of iWISDM lacks the capacity to probe these functions, it could potentially cover some of these additional capabilities by adding new operators like Count (to tally specific objects) or Relative (to identify properties like location in relation to other objects).
iWISDM provides an experimental environment for testing various continual learning (CL) algorithms, which address how artificial systems learn sequentially across a range of tasks. This platform can generate a variety of non-stationary tasks with increasing difficulty. Unlike most current multi-task learning models tested in fixed environments with relatively homogeneous task structures, iWISDM introduces dynamic task variations. For instance, in computer vision, CL algorithms typically classify an increasing number of image classes using datasets like MNIST (Lecun et al., 1998), CORe50 (Lomonaco and Maltoni, 2017), or similar datasets. Task-free continual learning approaches (Aljundi et al., 2019) exist but are still limited in generating a continual stream of tasks. In contrast, iWISDM allows users to generate a series of tasks with monotonically increasing or decreasing difficulty levels by adjusting complexity parameters in a fully controllable manner.
Achieving human-level intelligence requires models capable of out-of-distribution generalization, with compositionality being a promising solution to this challenge. Thus, it is important to evaluate whether a model can achieve compositional generalization. iWISDM’s graph structure allows for the manual construction of novel tasks based on existing task graphs, enabling systematic testing of both logical and temporal compositional generalization abilities. This makes iWISDM an ideal testbed for modularity-based CL algorithms, as it facilitates parallel connections between modules of models and tasks.
@@ Appendix A Appendix
A.1 Additional Figures
@@
Figure A2: An example of low complexity task generated by iWISDM.
Figure A3: An example of medium complexity task generated by iWISDM.
Figure A4: An example of high complexity task generated by iWISDM.
Figure A5: An example trial with distractors.
@@
Figure A7: Average accuracy of GPT-4v, xcomposer2, and MMICL across a varying number of delay frames.
@@
A.2 Model evaluation prompts
A.2.1 GPT-4V, Claude-3, & Gemini-pro Low/Medium Complexity (all properties + examples included)
@@
A.2.2 GPT-4V Single-Frame Complexity (location property + no examples included)
@@
A.2.3 GPT-4V, Claude-3, & Gemini-pro High Complexity (all properties + examples included)
@@
A.2.4 InternLM-XComposer2 & MMICL Low Complexity (all properties included + examples excluded)
@@
A.2.5 InternLM-XComposer2 High Complexity (all properties included + examples excluded)

### v3

@@ 4.2 Complexity
Using the AutoTask framework in iWISDM, we created three benchmarks corresponding to Low, Medium, and High complexity tasks. Low complexity tasks were restricted to contain exactly one logical joining operator (”and/or”), exclude Switch operators, and require only boolean actions. Medium complexity tasks had the same logical joining operator restriction as low complexity and boolean action requirements, with the complexity increase coming from adding a Switch operator. Finally, high complexity tasks have between one and two logical joining operators, an included Switch operator, and require boolean and object property action responses (i.e. ’… category of object 1?’, Answer: ’planes’). For further details refer to Figure A1. Additionally, since the benchmarks are generated based on a synthetic stimuli dataset, while all LMMs are trained with naturalistic stimuli, we aim to confirm whether LMMs can recognize rendered ShapeNet objects. Therefore, we developed two sets of simplistic single-frame tasks: location-only and category-only existence tasks (e.g. Is object 1 in the bottom left?). Evaluation on these two single-frame task sets provides an upper bound of LMMs’ performance.
The set of tested models was limited due to the scarcity of applicable models. Many popular open-source LMMs, such as LLaVa-1.5, were unable to perform the task simply due to their limited image sequence lengths. A minimum image sequence length of ten is needed to complete all complexity levels. We were able to properly evaluate two open-source models, InternLM-XComposer2-7b and MMICL-Instructblip-T5-xl. For samples of prompts used to evaluate each model see Appendix A.3.
@@ 4.3 Results
Figure 4a shows the accuracy of actions taken by the models plotted against the complexity level for each type of prompt. GPT-4V generally achieves the best model scores, with the largest performance gap on the low and high complexity tasks. However, relative to human performance these gaps are marginal at best. Broadly, MMICL was the worst-performing model. The expected inverse correlation between complexity and action accuracy was only captured clearly by the GPT-4V and Gemini-Pro-1.0 results.
@@
We also analyzed how each model performed on subsets of tasks where single or multiple object properties were involved. Figure 4 b-e shows the average accuracy of all models conditioned on task properties and complexity level. We did not include the object-identity subgroup in the high complexity plot of Figure 4e as non-boolean response types for object identity tasks are not feasible. The performances on high complexity tasks show a clear ranking between the abilities of GPT-4V and Gemini-pro models on tasks with diverse response options. For almost all models and complexities, location-only tasks posed the most difficulties. This finding confirms previous analyses of GPT-4V, which found that it often struggles to correctly recognize an object’s position within an image [^1] https://blog.roboflow.com/gpt-4v-object-detection (Majumdar et al., 2024).
To further investigate the response patterns of the LMMs we performed further analyses on a subset of the models seen in Appendix Figures A7, A8, A9, and A10. To determine the effects of delay frames on model performance a set of simple delayed-match-to-sample tasks were generated which only differed in difficulty by the number of delay frames. Appendix Figure A7 shows that for InternLM-XComposer2, MMICL, and GPT-4v, the addition of single or multiple delay frames in a task has little effect on task performance. Next, we looked into how the number of different boolean operators affected open-source model (InternLM-XComposer2 & MMICL) performance. The Appendix Figure A8 a-d results show that, generally, across all boolean operators an increase in their abundance within a task leads to worse model performance, which is expected. In Appendix Figure A9 we examined how the number of stimuli affected task performance across complexities. We expected to see that as the number of stimuli increases, the task performance would decrease. This was found to be the case for Low complexity tasks, which can be seen in Appendix Figure A9 a. However Medium and High complexity tasks displayed an inverse trend, as seen in Appendix Figure A10 b and c. Finally, we investigated the exact effect that different required response types had on accuracy for the High complexity benchmark. As Appendix Figure A10 shows, when tasks required a response that was a non-boolean type word, the models performed significantly worse.
@@ 5 Conclusion and future directions
We introduced iWISDM as a platform for validating multimodal models. As a benchmark, our primary focus is on assessing the ability of LMMs to follow instructions in visual-language decision-making tasks. We developed three benchmarks of incremental complexity and evaluated several LMMs and human performance. The gap between LMMs and humans indicates that LMMs still lack key abilities to solve instruction-following tasks. Through a detailed analysis of LMMs’ behaviour patterns, we identified diminished spatial recognition ability and decreased performance with increased task complexity.
We believe iWISDM would be an important benchmark that complements existing benchmarks which evaluate LMM capabilities in areas such as commonsense reasoning, numerical computation, or relational inferences (Fu et al., 2023). Although the current version of iWISDM lacks the capacity to probe these functions, it could potentially cover some of these additional capabilities by adding new operators like Count (to tally specific objects) or Relative (to identify properties like location in relation to other objects). We also believe iWISDM can serve as an important benchmark for continual learning algorithms evaluation. For detailed discussions, please refer to Appendix A.1.1.
@@ Appendix A Appendix
A.1 Additional Discussion
A.1.1 Extensibility of iWISDM
  • Visual inputs. iWISDM also allows any natural image set to be used as the stimulus set. In our core build, we use 2D projection of 3D object models from the ShapeNet dataset (Chang et al., 2015), where each stimulus is defined by a parameter vector consisting of category, identity, pose angle, location. We provide a template that allows users to seamlessly import alternative stimulus datasets.
A.1.2 iWISDM as a framework for Continual Learning
Additionally, iWISDM provides a framework for testing and comparing different approaches for continual learning (CL) and multi-task learning. iWISDM can be used to generate any number of tasks with quantifiable similarity which can in turn be used to test CL approaches on their capacity to sequentially learn tasks with different degrees of similarities without forgetting. Unlike most current multi-task learning models tested in fixed environments with relatively homogeneous task structures, iWISDM introduces dynamic task variations. For instance, in computer vision, CL algorithms are typically tested on classifying an increasing number of image classes using datasets such as MNIST (Deng, 2012), CORe50 (Lomonaco & Maltoni, 2017), or similar datasets. Task-free continual learning approaches (Aljundi et al., 2019) exist but are still limited in generating a continual stream of tasks. In contrast, iWISDM allows users to generate a series of tasks with monotonically increasing or decreasing difficulty levels by adjusting complexity parameters in a fully controllable manner.
Achieving human-level intelligence requires models capable of out-of-distribution generalization. Compositional generalization is an important ability of intelligent agents and thus, it is important to evaluate models on their ability to generalize to compositionally generated scenarios. iWISDM’s graph structure allows for the manual construction of novel tasks based on existing task graphs, enabling systematic testing of both logical and temporal compositional generalization abilities. This makes iWISDM an ideal test bed for modularity-based CL algorithms, as it facilitates parallel connections between modules of models and tasks.
A.2 Additional Figures
@@
Figure A2: An example of low complexity task generated by iWISDM. Instruction: ”observe object 1, observe object 2, delay, observe object 3, observe 4, location of object 3 equals location of object 2 or location of object 1 equals location of object 4?”
Figure A3: An example of medium complexity task generated by iWISDM.Instruction: ”observe object 1, observe object 2, observe object 3, observe 4, observe object 5, observe object 6, observe object 7, observe 8, if location of object 7 equals location of object 2 or location of object 8 equals location of object 4, then location of object 3 equals location of object 1? else location of object 6 equals location of object 5?”
Figure A4: An example of high complexity task generated by iWISDM. Instruction: ”observe object 1, observe object 2, delay, observe object 3, observe 4, observe object 5, delay, observe object 6, observe object 7, if location of object 3 not equals top right and location of object 2 equals location of object 5, then location of object 7 equals top left and location of object 6 equals location of object 4? else location of object 1?”
Figure A5: An example trial with distractors. Instruction: ”observe object 1, observe object 2, category of object 2 not equals category of object 1? observe object 3,observe object 4 with location:, top left, category of object 4 equals couches or identity of object 3 equals identity of object 1? delay, observe object 5, observe object 6, delay, observe object 7, identity of object 7 equals identity of object 6 and identity of object 4 equals identity of object 3?”
@@
Figure A7: Additional Analyses of Open-Source model results (xcomposer2 & MMICL)
@@
A.3 Model evaluation prompts
A.3.1 GPT-4V, Claude-3, & Gemini-pro Low/Medium Complexity (all properties + examples included)
In this task we will show you a series of frame images. Each frame will either be blank (delay frame) or contain a 3D object.The objects within the task will ALWAYS be from one of 8 categories: benches, boats, cars, chairs, couches, lighting, planes, and tables. For each of these 8 categories, there are 8 unique objects that could be used in the task. Any object which is sampled will be displayed as an image taken from a random viewing angle. The objects will be placed in one of four locations: top left, top right, bottom left, and bottom right.
@@ A.3.1 GPT-4V, Claude-3, & Gemini-pro Low/Medium Complexity (all properties + examples included)
A.3.2 GPT-4V Single-Frame Complexity (location property + no examples included)
@@
A.3.3 GPT-4V, Claude-3, & Gemini-pro High Complexity (all properties + examples included)
@@
A.3.4 InternLM-XComposer2 & MMICL Low Complexity (all properties included + examples excluded)
@@
A.3.5 InternLM-XComposer2 High Complexity (all properties included + examples excluded)